# Nonlocal electrodynamics and the penetration depth of superconducting Sr2RuO4

### v1

@@ I Introduction
Obtaining a detailed understanding of the superconducting state in Sr2RuO4 remains an important outstanding problem [1, 2, 3]. Experimentally, the quest is complicated due to challenging material properties of Sr2RuO4 and the low energy scale of the superconducting phase. Theoretically, Sr2RuO4 provides an important testbed for modeling of unconventional superconductivity. More specifically, the well-characterized electronic properties of the normal state offers a rather unique opportunity to test various electronic fluctuation-based mechanisms for pairing against detailed measurements of the superconducting gap structure [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
The magnetic penetration depth, $\lambda$, can elucidate nodal features of the gap, as revealed through its dependency on temperature and disorder [24, 25, 26, 27]. Fully-gapped superconductors exhibit exponential temperature dependence at low temperatures, whereas nodal gaps display power law dependence. In Sr2RuO4, measurements of the change in penetration depth, $\Delta\lambda(T)$, via a tunnel diode oscillator method yielded $\Delta\lambda(T)\sim T^{2}$, signalling the existence of nodal quasiparticle excitations [28]. Renewed measurements of $\Delta\lambda(T)$ in Sr2RuO4 high-purity spherical single crystals by applying scanning SQUID microscopy [29] and ac-susceptibility measurements [30] have confirmed this behavior and highlighted the importance of nonlocal Meissner screening. Indeed, as initially demonstrated by Kosztin and Leggett, nonlocal effects may change the low temperature dependence of the penetration depth, e.g., from linear to quadratic in $d_{x^{2}-y^{2}}$-wave superconductors, such as the cuprates [31, 32]. The relevance of nonlocal effects in an extended temperature regime is controlled by the zero-temperature Ginzburg–Landau parameter, i.e., the ratio of the penetration depth to the coherence length. In Sr2RuO4 this ratio is $\kappa_{0}\approx 1.92$ [30], placing this material unusually close to the Pippard limit compared to most clean, unconventional superconductors. Finally, we note that the penetration depth was also recently extracted from SQUID susceptometry in thin Sr2RuO4 films, again yielding $\Delta\lambda(T)\sim T^{2}$ behavior at low $T$ [33]. There, however, the origin of the $T^{2}$ dependence was interpreted in terms of disorder scattering similar to disordered $d_{x^{2}-y^{2}}$-wave superconductors [34].
@@ II Penetration depth
Kosztin and Leggett showed that thermal quasiparticles in a clean, type-II, nodal $d$-wave superconductor are responsible for the behavior $\Delta\lambda\sim T^{2}$ for $T<T^{\ast}$ with $T^{\ast}=\Delta_{0}/\kappa_{0}$, where $\Delta_{0}$ is gap scale and $\kappa_{0}=\lambda_{0}/\xi_{0}$ ($\lambda_{0}$ and $\xi_{0}$ is the penetration depth and the coherence length, respectively) the zero-temperature Ginzburg–Landau parameter [31]. The key ingredient here are the nonlocal effects (diverging coherence length) experienced by the Cooper pairs formed by momenta near the nodal points, see Fig. 1. This causes the penetration depth increase to acquire an additional factor of $T$ (on top of the $2k_{B}T\ln{2}/\Delta_{0}$ local contribution) coming from the inverse thermal de Broglie wavelength at low temperatures.
@@ II.1 Nonlocal electrodynamics
where $\tilde{q}=q\lambda_{0}$. We emphasize that this is an exact expression that includes a correction responsible for an upturn in $\Delta\lambda$ at higher $T$ as compared to Ref. 31. The dimensionless kernel can be computed by means of standard Green’s function methods, assuming linear response in the Meissner state and solving the relevant Maxwell equation in the superconductor [38]. In the Matsubara representation the kernel can be expressed as [38, 31]
@@
where the fermionic Matsubara frequencies are given by $\omega_{n}=\pi k_{B}T(2n+1)$, $\hat{p}_{\parallel}=\cos(\theta)$ is the Fermi surface momentum projected onto the boundary ($\theta$ is the polar angle of $\boldsymbol{p}$), and the parameter $\alpha=\pi\Delta_{0}\tilde{q}\sin(\theta)/(2\kappa_{0})$ is the projected magnetic field penetration and is responsible for the nonlocal effects. Here, we employed the BCS expression for the coherence length, such that $\lambda_{0}=v_{F}\kappa_{0}/(\pi\Delta_{0})$. Finally, $\Delta_{\boldsymbol{p}}$ is the (temperature-dependent) order parameter, and the (dimensionless) Fermi surface average is evaluated as $\langle h\rangle_{\mathrm{FS}}\equiv(\bar{v}_{F}/\lvert S_{F}\rvert)\int_{S_{F}}\mathrm{d}\hat{\boldsymbol{k}}~{}h(\hat{\boldsymbol{k}})/v_{F}(\hat{\boldsymbol{k}})$ where $v_{F}(\hat{\boldsymbol{k}})$ is the Fermi velocity and $\lvert S_{F}\rvert$ the Fermi surface area. It can be noted that the zero-temperature kernel, $\tilde{K}(\tilde{q};0)$, has a simple closed-form expression, as stated in Appendix A.
@@ II.2 The node approximation
There are at least two reasons why there are few candidate materials for observing $\Delta\lambda\sim T^{2}$ from the above mechanism. First, disorder is known to cause a saturation of $\Delta\lambda$ below a characteristic temperature set by the scattering rate [34], which will effectively blur the observation of a quadratic temperature dependence. Second, the most well-characterized $d$-wave superconductors are strong type-II, $\kappa_{0}\gg 1$, suppressing $T^{\ast}$ to a tiny fraction of the gap. In both of these respects Sr2RuO4 poses as a counterexample, since it has accumulated strong evidence of being both nodal [42, 43], marginal type-II [30], and is known to produce clean and disorder-sensitive crystals [44]. Other candidate superconductors possibly relevant to a low-$T$ penetration depth of $\Delta\lambda\sim T^{2}$ caused by nonlocal electrodynamics include cuprates [31], KFe2As2 [45], and the heavy-fermion material CeCoIn5 [46, 24]. Below we focus on the case of Sr2RuO4.
@@ III.2 Numerical evaluation of the penetration depth
First, we calculate the kernel difference $-\delta\tilde{K}(\tilde{q};T)$ at two temperatures using both the node approximation of Eq. (3) and the full Matsubara sum of Eq. (2). For illustration we employ the simplest gap structure consistent with B1g symmetry, i.e., $\Delta_{0}\cos(2\theta)$ on all bands with $\Delta_{0}=0.35$ meV on bands $\alpha$ and $\beta$, and half the magnitude on band $\gamma$. These values are motivated by STM experiments [36, 29], and the gap magnitude on $\gamma$ was reduced to match the experimental penetration depth slope at low $T$. We otherwise fixed $\kappa_{0}=1.92$ and $T_{c}=1.5$ K, in agreement with experiments [30]. The resulting kernel differences are shown in Fig. 3, and the order parameter and penetration depth are shown in Fig. 4(a) and (c).
Comparing the penetration depth calculations with the experimental data reveals that the temperature window in which the node approximation matches with the realistic modelling is roughly $T/T_{c}\lesssim 0.2$. The primary reason for this is the overestimated slope of $\lvert\Delta(\theta)\rvert$ when linearizing the gap, causing the node approximation to, in this case, monotonically underestimate the penetration depth. Additionally, the $T$ dependence of the gap and the correction posed by the denominator of Eq. (1) both contribute with an upturn in $\Delta\lambda(T)$ close to the transition temperature $T_{c}$, approximately consistent with the experimental data. The realistic modelling shows that a simple $d_{x^{2}-y^{2}}$-wave gap is sufficient to explain the experimental data, albeit with a slight discrepancy at the highest $T$.
@@ Appendix A Generalized Kosztin–Leggett theory
where $\alpha_{j}=\tilde{q}v_{F,j}\sin(\theta_{j})/(2\lambda_{0})$, $\tilde{f}(x)\equiv(1+\exp(x))^{-1}$ is the Fermi function with dimensionless argument, and where the sum runs over distinct nodes in the Brillouin zone (possibly distributed across multiple bands). This result is a multiband generalization of the node approximation proposed by Kosztin and Leggett, which is a low-temperature approximation that is valid in the temperature window in which the order parameter can be reasonably approximated by a linear function of the angle deviation from the node.
@@
$-\delta\tilde{K}(\tilde{q};T)$ $=\frac{2k_{B}T\ln{2}}{\Delta_{0}}$ (16)
×$\times\left[1-\frac{1}{\ln{2}}\int_{0}^{\pi\frac{\sqrt{2}z}{4}}\mathrm{d}x~{}\tilde{f}(x)\sqrt{1-\frac{8x^{2}}{\pi^{2}z^{2}}}\right],$
where $z\equiv\frac{\tilde{q}}{T/T^{\ast}}$, and $T^{\ast}\equiv\frac{\Delta_{0}}{\kappa_{0}}$, where $\kappa_{0}$ is the zero-temperature Ginzburg–Landau parameter, and the prefactor is recognized as $-\delta\tilde{K}(0;T)=2k_{B}T\ln{2}/\Delta_{0}$.

### v2

@@ I Introduction
Obtaining a detailed understanding of the superconducting state in Sr2RuO4 remains an important outstanding problem [1, 2, 3]. Experimentally, the quest is complicated due to challenging material properties of Sr2RuO4 and the low energy scale of the superconducting phase. Theoretically, Sr2RuO4 provides an important testbed for modeling of unconventional superconductivity. More specifically, the well-characterized electronic properties of the normal state offers a rather unique opportunity to test various electronic fluctuation-based mechanisms for pairing against detailed investigations of the superconducting gap structure [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
The magnetic penetration depth, $\lambda$, can elucidate nodal features of the gap, as revealed through its dependency on temperature and disorder [24, 25, 26, 27]. Fully-gapped superconductors exhibit exponential temperature dependence at low temperatures, whereas nodal gaps display power law dependence. In the case of a simple d-wave superconductor, deviations linear in the temperature $T$ are expected and a canonical interpretation of a $T^{2}$ dependence in d-wave superconductors is impurities. In Sr2RuO4, measurements of the change in penetration depth, $\Delta\lambda(T)\equiv\lambda(T)-\lambda(0)$, via a tunnel diode oscillator method also yielded $\Delta\lambda(T)\sim T^{2}$, signalling the existence of nodal quasiparticle excitations [28]. Renewed measurements of $\Delta\lambda(T)$ in Sr2RuO4 high-purity spherical single crystals using scanning SQUID microscopy [29] and ac-susceptibility measurements [30] have confirmed this behavior, excluded interpretation in terms of impurity effects and highlighted the importance of nonlocal Meissner screening. Indeed, as initially demonstrated by Kosztin and Leggett, nonlocal effects change the low temperature dependence of the penetration depth, e.g., from linear to quadratic in $d_{x^{2}-y^{2}}$-wave superconductors, such as the cuprates [31, 32]. The relevance of nonlocal effects in an extended temperature regime is controlled by the zero-temperature Ginzburg–Landau parameter, i.e., the ratio of the penetration depth to the coherence length. In Sr2RuO4 this ratio is $\kappa_{0}\approx 1.92$ [30], placing this material unusually close to the Pippard limit compared to most clean, unconventional superconductors. Finally, we note that the penetration depth was also recently extracted from SQUID susceptometry in thin Sr2RuO4 films, again yielding $\Delta\lambda(T)\sim T^{2}$ behavior at low $T$ [33]. There, however, the origin of the $T^{2}$ dependence was interpreted in terms of disorder scattering similar to disordered $d_{x^{2}-y^{2}}$-wave superconductors [34].
@@ II Penetration depth
Kosztin and Leggett showed that thermal quasiparticles in a clean, type-II, nodal $d$-wave superconductor are responsible for the behavior $\Delta\lambda\sim T^{2}$ for $T<T^{\ast}$ with $T^{\ast}=\Delta_{0}/\kappa_{0}$, where $\Delta_{0}$ is gap scale and $\kappa_{0}=\lambda_{0}/\xi_{0}$ ($\lambda_{0}$ and $\xi_{0}$ is the penetration depth and the coherence length, respectively) the zero-temperature Ginzburg–Landau parameter [31]. The key ingredient here are the nonlocal effects (diverging coherence length) experienced by the Cooper pairs formed by momenta near the nodal points, see Fig. 1. For such Cooper pairs, it is important to include the spatial variation of the vector potential.  [^1] We note that a different kind of nonlocal Meissner effect based on anisotropies within a multiband London model was studied in Ref. [53]. This causes the penetration depth increase to acquire an additional factor of $T$ (on top of the $2k_{B}T\ln{2}/\Delta_{0}$ local contribution) coming from the inverse thermal de Broglie wavelength at low temperatures.
@@ II.1 Nonlocal electrodynamics
where $\tilde{q}=q\lambda_{0}$, see Appendix A. We emphasize that this is an exact expression that includes a correction responsible for an upturn in $\Delta\lambda$ at higher $T$ as compared to Ref. 31. The dimensionless kernel can be computed by means of standard Green’s function methods, assuming linear response in the Meissner state and solving the relevant Maxwell equation in the superconductor [39]. In the Matsubara representation the kernel can be expressed as [39, 31]
@@
where the fermionic Matsubara frequencies are given by $\omega_{n}=\pi k_{B}T(2n+1)$, $\hat{p}_{\parallel}=\cos(\theta)$ is the Fermi surface momentum projected onto the boundary ($\theta$ is the polar angle of $\boldsymbol{p}$), and the parameter $\alpha=\pi\Delta_{0}\tilde{q}\sin(\theta)/(2\kappa_{0})$ is the projected magnetic field penetration and is responsible for the nonlocal effects. Here, we employed the BCS expression for the coherence length, such that $\lambda_{0}=v_{F}\kappa_{0}/(\pi\Delta_{0})$. Finally, $\Delta_{\boldsymbol{p}}$ is the (temperature-dependent) order parameter, and the (dimensionless) Fermi surface average is evaluated as $\langle h\rangle_{\mathrm{FS}}\equiv(\bar{v}_{F}/\lvert S_{F}\rvert)\int_{S_{F}}\mathrm{d}\hat{\boldsymbol{k}}~{}h(\hat{\boldsymbol{k}})/v_{F}(\hat{\boldsymbol{k}})$ (see Eq. (10)) where $v_{F}(\hat{\boldsymbol{k}})$ is the Fermi velocity and $\lvert S_{F}\rvert$ the Fermi surface area. It can be noted that the zero-temperature kernel, $\tilde{K}(\tilde{q};0)$, has a simple closed-form expression, as stated in Appendix A.
@@ II.2 The node approximation
There are at least two reasons why there are few candidate materials for observing $\Delta\lambda\sim T^{2}$ from the above mechanism. First, disorder is known to cause a saturation of $\Delta\lambda$ below a characteristic temperature set by the scattering rate [34], which will effectively blur the observation of a quadratic temperature dependence. Second, the most well-characterized $d$-wave superconductors are strong type-II, $\kappa_{0}\gg 1$, suppressing $T^{\ast}$ to a tiny fraction of the gap. In both of these respects Sr2RuO4 presents a counter-example: there is strong evidence that it is both nodal [43, 44] and only marginally type-II [30], and it can be prepared to exceptionally high crystal quality [45]. Other candidate superconductors possibly relevant to a low-$T$ penetration depth of $\Delta\lambda\sim T^{2}$ caused by nonlocal electrodynamics include cuprates [31], KFe2As2 [46], and the heavy-fermion material CeCoIn5 [47, 24, 48]. Below we focus on the case of Sr2RuO4.
@@ III.2 Numerical evaluation of the penetration depth
First, we calculate the kernel difference $-\delta\tilde{K}(\tilde{q};T)$ at two temperatures using both the node approximation of Eq. (3) and the full Matsubara sum of Eq. (2). For illustration we employ the simplest gap structure consistent with B1g symmetry, i.e., $\Delta_{0}\cos(2\theta)$ on all bands with $\Delta_{0}=0.35$ meV on bands $\alpha$ and $\beta$, and half the magnitude on band $\gamma$. These values are motivated by STM experiments [37, 29], and the gap magnitude on $\gamma$ was reduced to match the experimental penetration depth slope at low $T$. We otherwise fixed $\kappa_{0}=1.92$ and $T_{c}=1.5$ K, in agreement with experiments [30]. The resulting kernel differences $-\delta\tilde{K}(\tilde{q};T)$ are shown in Fig. 3, and the order parameter and penetration depth are shown in Fig. 4(a) and (c).
Comparing the penetration depth calculations with the experimental data reveals that the temperature window in which the node approximation matches with the realistic modelling is roughly $T/T_{c}\lesssim 0.2$. The primary reason for this is that the node approximation based on the slope $v_{\Delta}$ overestimates the value of the order parameter $\lvert\Delta(\theta)\rvert$ when linearizing the gap. For a given temperature $T$, there are fewer quasiparticle states available in the node approximation compared to the full dependence of $\Delta(\theta)$, causing the node approximation to, in this case, monotonically underestimate the penetration depth. Additionally, the $T$ dependence of the gap and the correction posed by the denominator of Eq. (1) both contribute with an upturn in $\Delta\lambda(T)$ close to the transition temperature $T_{c}$, approximately consistent with the experimental data. The realistic modelling shows that a simple $d_{x^{2}-y^{2}}$-wave gap is sufficient to explain the experimental data, albeit with a slight discrepancy at the highest $T$.
@@
In the above analysis we have focused on $d$-wave ordering since it appears as a strong contender for the leading superconducting order relevant for Sr2RuO4. We note, however, that purely based on the penetration depth data, only the nodal structure of the gap function matters. Thus, one may also obtain equally good fits to the experimental data shown in Fig. 4 (c,d) from, for example, a nodal $s$-wave order parameter. Finally, fully gapped order parameters such as the (earlier) proposed triplet state of type $p_{x}+ip_{y}$ is incapable of explaining the penetration depth data of Sr2RuO4.
@@ Appendix A Generalized Kosztin–Leggett theory
where $\alpha_{j}=\tilde{q}v_{F,j}\sin(\theta_{j})/(2\lambda_{0})$, $\tilde{f}(x)\equiv(1+\exp(x))^{-1}$ is the Fermi function with dimensionless argument, and where the sum runs over distinct nodes in the Brillouin zone (possibly distributed across multiple bands). This result is a multiband generalization of the node approximation proposed by Kosztin and Leggett. This low-temperature approximation is valid in the temperature window in which the order parameter can be reasonably approximated by a linear function of the angle deviation from the node.
@@
$-\delta\tilde{K}(\tilde{q};T)$ $=\frac{2k_{B}T\ln{2}}{\Delta_{0}}\left[1-\frac{1}{\ln{2}}\int_{0}^{\pi\frac{\sqrt{2}z}{4}}\mathrm{d}x~{}\tilde{f}(x)\sqrt{1-\frac{8x^{2}}{\pi^{2}z^{2}}}\right],$ (16)
where $z\equiv\frac{\tilde{q}}{T/T^{\ast}}$, and $T^{\ast}\equiv\frac{\Delta_{0}}{\kappa_{0}}$. Here $\kappa_{0}$ is the zero-temperature Ginzburg–Landau parameter, and the prefactor is recognized as $-\delta\tilde{K}(0;T)=2k_{B}T\ln{2}/\Delta_{0}$.